# Renormalisation for Reaction-Diffusion Systems with Non-Local Interactions

### v1

@@ 2.5 Mean Field Behaviour
So the end result is just that seen for local interaction. Now the steady state of the annihilation Model I is population extinction (or a lone survivor), and it is the time decay towards this state that is of interest. In particular, it is known [Lee1994] for local interaction and large times (without mean field approximation), that below critical dimension $d_{c}=2$, the population decay is of order $\mathcal{O}(t^{-d/2})$, whereas above this dimension, it is just of order $\mathcal{O}(t^{-1})$, as seen in the mean field solution in Eq. (7).
@@
Thus the same effect results as seen in Model I; the tree structure contains zero (spatial) momentum, interaction terms such as $R(k=0)=R\hat{R}(k=0)=R$ simplify, and the mean field equation observed for local interaction [Janssen1999] results.
In [Janssen1999], it is known (without mean field approximation) for local interaction that the position where positive density emerges is shifted from $\tau=0$. The scaling also differs, if $\bar{\tau}$ is the distance above this shifted point, it takes the form $X\propto\bar{\tau}^{\beta}$ for non-unit constant $\beta=1-\frac{\epsilon}{6}+\mathcal{O}(\epsilon^{2})$ (where $\epsilon=4-d$).
@@ 3 Renormalisation for Annihilation Model
Renormalisation methods can be used to deal with divergences seen in perturbative expansions, and also to examine critical exponents in phase transitions [Bellac1992, Binney1992, Collins1984, Lee1994, Talagrand2022, Tauber2014, Vasil2004].
In this section, the renormalisation techniques seen in [Lee1994] are first adapted to non-local interaction for annihilation Model I (the addition of branching, spontaneous birth and death in Model II is considered in the Section 4). Thus $Q_{pq}=M_{p}=B_{p}=0$ throughout. In this section, the nature of UV and IR divergences are first highlighted in more detail. It is then shown how solutions to the Callan-Symanzyk equation found in [Lee1994] can be derived by a simple scaling argument without need of the equation itself. This is then used to show that in the asymptotic regime, non-local coupling becomes local and the same asymptotics apply.
@@ 3.1 Ultra Violet Regulation and Infra Red Divergence
$I_{2}$ $=$ $R^{2}\int_{t>t_{2}>t_{1}>0}\fint_{k}e^{-\frac{k^{2}}{2\lambda^{2}}}e^{-2Dk^{2}(t_{2}-t_{1})},$
$=$ $\frac{R^{2}}{(8\pi)^{\frac{d}{2}}}\left\{\frac{1}{(2D)^{2}(2-d)(4-d)}\left[(\lambda^{-2}+4Dt)^{-\frac{d}{2}+2}-\lambda^{d-4}\right]-\frac{\lambda^{d-2}t}{2D(2-d)}\right\},$
@@
Now the second line reduces to $\frac{R^{2}t^{-\frac{d}{2}+2}}{(8\pi D)^{\frac{d}{2}}(2-d)(4-d)}$ as $\lambda\rightarrow\infty$. This expression can be found in [Tauber2005] for the local interaction. There it is noted that the effective (dimensionless) coupling $I_{2}/I_{1}$ is proportional to $t^{-\frac{d}{2}+1}$ which for $d>2$ diverges for small time (this is essentially the UV divergence). For $d<2$ this term also diverges in the (IR) regime of large $t$. Now, for finite $\lambda$, and $t<\frac{1}{4D\lambda^{2}}$ sufficiently small, a series expansion in $t$ can be found which cancels the removable d2d-2 and d4d-4 singularities and gives a convergent series for any dimension and the UV divergence is naturally regulated. For large $t$ and $d>2$ the ratio $I_{2}/I_{1}$ is still well behaved. However, for large $t$ and $d<2$ we find (from the second line above) that we still have an IR divergence to deal with.
If the screened Poisson profile is used instead, the following integral is obtained. Note that the two interaction propagators (which just become $R^{2}$ in the local model) each provide an extra quadratic factor in the denominator and primitive divergence estimates suggest convergence for $d<6$. If Schwinger parametrisation [Collins1984, Smirnov2006, Weinzierl2022] is initially used, followed by the $t$ integration (using convolution properties of Laplace transforms), and finally the the momentum integration over $k$ performed, we obtain,
@@
$=$ $\frac{R^{2}}{(4\pi)^{\frac{d}{2}}}\int_{0}^{\infty}\mbox{d}x\hskip 1.42262ptxe^{-x}\left[\frac{\left(\frac{x}{\lambda^{2}}+2Dt\right)^{-\frac{d}{2}+2}-\left(\frac{x}{\lambda^{2}}\right)^{-\frac{d}{2}+2}}{D^{2}(2-d)(4-d)}-\frac{(\frac{x}{\lambda^{2}})^{-\frac{d}{2}+1}t}{D(2-d)}\right]$
$=$ $\frac{R^{2}\lambda^{d-4}\left[e^{2Dt\lambda^{2}}\Gamma\left(-\frac{d}{2}+4,2Dt\lambda^{2}\right)-2Dte^{2Dt\lambda^{2}}\Gamma\left(-\frac{d}{2}+3,2Dt\lambda^{2}\right)-(1+(4-d)Dt)\Gamma(-\frac{d}{2}+4)\right]}{(4\pi)^{\frac{d}{2}}D^{2}(d-2)(d-4)}.$
@@
In the second line above the transform $y=\frac{x}{\lambda^{2}}+2Dt$ is used in the first term and then upper incomplete gamma functions can be used. Note both factors d2d-2 and d4d-4 are removable singularities, which can be seen by L’Hopital’s rule upon the second line. For finite $t$ the incomplete upper gamma functions are entire and contain no divergences. The only divergence occurs in the final term for $d=8,10,12,\dots$ and is otherwise regulated. However, as $t$ approaches zero, the integral becomes divergent at $d=6$ as the incomplete gamma function becomes complete (with pole), matching the primitive dimension of divergence. Thus the UV regulation is found to be weaker than for the normal interaction, which has an exponentially decaying tail and the integral is convergent in any dimension. For $d<2$ we again find IR behaviour for large $t$.
For the spherical interaction, these calculations are somewhat intractable, although the slow decay of the asymptotic form $R(k)\propto k^{-\frac{d}{2}}J_{\frac{d}{2}}(\frac{k}{\lambda})\sim k^{-\frac{d}{2}}\sqrt{\frac{2\lambda}{\pi k}}\cos(\frac{k}{\lambda}-(d+1)\frac{\pi}{4})$ suggests that the regulation is weaker than either the normal or screened Poisson interaction for $d\approx 2$. This may reflect the highly localised nature of the spherical interaction.
@@
Thus we find that when the precision parameter is finite, the non-local interactions act as a natural form of regulation (in the renormalisation sense), suggesting control of UV divergences is possible provided the interaction function tails off sufficiently quickly in momentum representation. Note, however, that as the precision parameter $\lambda$ for any of these non-local models gets larger, the regulation weakens as the interaction tends toward local behaviour and UV divergences start to emerge. Even for finite precision though, IR behaviour is still present for large $t$, for which we turn to rescaling methods that underlie the renormalisation group.
@@ 3.2 Rescaling
Here it is shown that the renormalisation group calculation in [Lee1994] can also be found via a simple rescaling of the action in Eq. (4). Specifically, if we perform the following space-time-field transforms,
@@
Without renormalisation, this form of rescaling is not too involved. For example, it is a simple exercise to show the mean field solution in Eq. (7) satisfies this property. However, once renormalization is set at some scale, establishing this form of invariance is less straightforward. This expression is also somewhat akin to the solution to the Callan-Symanzyk equations seen in [Lee1994]. The main differences are firstly, there is rescaling of the extra parameter $\lambda$, and secondly, that it is in terms of the original interaction parameter $R$ rather than a renormalised version. However, we can recover the Callan-Symanzyk solution via the simple scaling above (without solving or needing the Callan-Symanzyk PDE) as follows.
In the case of local interactions, the following vertex function is obtained in momentum-Laplace space (see [Lee1994] for details),
@@
The constant $g^{*}=\frac{(4\pi)^{1-\frac{\epsilon}{2}}}{\Gamma(\frac{\epsilon}{2})}$ is of order $\mathcal{O}(\epsilon)$. This is just the expression given in [Lee1994] (which considers the more general (but local) $K$ particle process $KA\rightarrow\phi$) restricted to the case $K=2$. Next we select a fixed momentum scale $\kappa$ by defining $\kappa^{2}=s+p^{2}/2$ for some $s$ and $p$. In [Lee1994] and other applications it is convenient to pick zero momentum $p=0$ (and so $s=\kappa^{2}$), although the specific choice is immaterial and somewhat arbitrary. This is used to define a renormalised parameter $R_{r}$,
@@
This is precisely the form seen in [Lee1994] as part of the Callan-Symanzik solution. In particular, for $d<d_{c}=2$ we find that as $t\rightarrow\infty$ (so $\gamma\rightarrow 0$), the limit $g_{r}^{\prime}\rightarrow g^{*}$ arises, whereas for $d>d_{c}$ the limit $g_{r}^{\prime}\rightarrow 0$ occurs. These are precisely the fixed points of the Callan-Symanzyk equations in [Lee1994]. Note also that the relationship between $g_{r}$ and $g_{r}^{\prime}$ as we shift between $t$ and $t^{\prime}$ is independent of the choice $\kappa$, which does nothing other than set the scale and initialise the value $g_{r}$. Finally, we note that if the $R$ dependency is replaced with the dimensionless parameter $g_{r}$ (and we write $g_{r}^{\prime}=g_{r}(\gamma)$), the rescaling law in [Lee1994] is found, that is, the full solution to the Callan-Symanzyk equation for the density;
@@
Note that from $\gamma^{2}=\frac{t^{\prime}}{t}$ we find that for large time, the precision parameter, scaling as $\lambda\gamma^{-1}$, becomes large, and the interaction profiles become local. This ultimately explains why the asymptotics of non-local systems are the same as local systems, although the cross-over time scale of such behaviours is a natural question to consider.
@@ 3.3 The Response Functional
The response functional is the standard propagator running from time $t_{1}$ to $t_{2}$ with any number of tree branches attached, where each tree runs from time 0 to the time of attachment somewhere in $t_{1},t_{2}]$. The first few corresponding diagrams can be seen in Fig. 5 (represented as spirals). The external momenta running along attached trees are taken to be zero (this essentially means the initial densities $n_{0}$ at the end of trees are spatially uniform). The momenta running along the response functional, $k$ (which may be part of a loop in a larger diagram, for example), can be non-zero. Note that each each time point of attachment, we have two possible nodes to attach the tree branch to. This results in two possible choices for the interaction propagator, either $R(0)$ or $R(k)$, leading to the following expression.
@@
$=e^{-Dk^{2}(t_{2}-t_{1})}\left(\frac{1+Rn_{0}t_{1}}{1+Rn_{0}t_{2}}\right)^{\frac{R(0)+R(k)}{R}}=e^{-Dk^{2}(t_{2}-t_{1})}\left(\frac{1+Rn_{0}t_{1}}{1+Rn_{0}t_{2}}\right)^{1+\hat{R}(k)}$

### v2

@@ 2.5 Mean Field Behaviour
So the end result is just that seen for local interaction. Now the steady state of the annihilation Model I is population extinction (or a lone survivor), and it is the time decay towards this state that is of interest. In particular, it is known [23] for local interaction and large times (without mean field approximation), that below critical dimension $d_{c}=2$, the population decay is of order $\mathcal{O}(t^{-d/2})$, whereas above this dimension, it is just of order $\mathcal{O}(t^{-1})$, as seen in the mean field solution in Eq. (7).
@@
Thus the same effect results as seen in Model I; the tree structure contains zero (spatial) momentum, interaction terms such as $R(k=0)=R\hat{R}(k=0)=R$ simplify, and the mean field equation observed for local interaction [19] results.
In [19], it is known (without mean field approximation) for local interaction that the position where positive density emerges is shifted from $\tau=0$. The scaling also differs, if $\bar{\tau}$ is the distance above this shifted point, it takes the form $X\propto\bar{\tau}^{\beta}$ for non-unit constant $\beta=1-\frac{\epsilon}{6}+\mathcal{O}(\epsilon^{2})$ (where $\epsilon=4-d$).
@@ 3 Renormalisation for Annihilation Model
Renormalisation methods can be used to deal with divergences seen in perturbative expansions, and also to examine critical exponents in phase transitions [3, 4, 7, 23, 30, 32, 35].
In this section, the renormalisation techniques seen in [23] are first adapted to non-local interaction for annihilation Model I (the addition of branching, spontaneous birth and death in Model II is considered in the Section 4). Thus $Q_{pq}=M_{p}=B_{p}=0$ throughout. In this section, the nature of UV and IR divergences are first highlighted in more detail. It is then shown how solutions to the Callan-Symanzyk equation found in [23] can be derived by a simple scaling argument without need of the equation itself. This is then used to show that in the asymptotic regime, non-local coupling becomes local and the same asymptotics apply.
@@ 3.1 Ultra Violet Regulation and Infra Red Divergence
$I_{2}$ $=$ $R^{2}\int_{t>t_{2}>t_{1}>0}\fint_{k}e^{-\frac{k^{2}}{2\lambda^{2}}}e^{-2Dk^{2}(t_{2}-t_{1})}$
$=$ $\frac{R^{2}}{(8\pi)^{\frac{d}{2}}}\left\{\frac{1}{(2D)^{2}(2-d)(4-d)}\left[(\lambda^{-2}+4Dt)^{-\frac{d}{2}+2}-\lambda^{d-4}\right]-\frac{\lambda^{d-2}t}{2D(2-d)}\right\}$
@@
Now the second line reduces to $\frac{R^{2}t^{-\frac{d}{2}+2}}{(8\pi D)^{\frac{d}{2}}(2-d)(4-d)}$ as $\lambda\rightarrow\infty$. This expression can be found in [31] for the local interaction. There it is noted that the effective (dimensionless) coupling $I_{2}/I_{1}$ is proportional to $t^{-\frac{d}{2}+1}$ which diverges for $d>2$ and small time (this is essentially the UV divergence). For $d<2$ this term also diverges in the (IR) regime of large $t$. Now, for finite $\lambda$, and $t<\frac{1}{4D\lambda^{2}}$ sufficiently small, a series expansion in $t$ can be found which cancels the removable d2d-2 and d4d-4 singularities and gives a convergent series for any dimension and the UV divergence is naturally regulated. For large $t$ and $d>2$ the ratio $I_{2}/I_{1}$ is still well behaved. However, for large $t$ and $d<2$ we find (from the second line above) that we still have an IR divergence to deal with.
Note that these observations also point towards a cross-over timescale between local and non-local features. More specifically, we have already observed that as the precision parameter $\lambda\rightarrow\infty$, the expression for $I_{2}$ given by the local model is recovered, as expected. Furthermore, we see that from the condition $t<\frac{1}{4D\lambda^{2}}$, as $\lambda$ increases towards this limit, the range of $t$ values for which UV regulation exists decreases, and the model naturally becomes more local.
If the screened Poisson profile is used instead, the following integral is obtained for $I_{2}$. Note that the two interaction propagators (which just become $R^{2}$ in the local model) each provide an extra quadratic factor in the denominator and primitive divergence estimates suggest convergence for $d<6$. If Schwinger parametrisation [7, 28, 37] is initially used, followed by the $t$ integration (using convolution properties of Laplace transforms), and finally the the momentum integration over $k$ performed, we obtain,
@@
$=$ $\frac{R^{2}}{(4\pi)^{\frac{d}{2}}}\int_{0}^{\infty}\mbox{d}x\hskip 1.42262ptxe^{-x}\left[\frac{\left(\frac{x}{\lambda^{2}}+2Dt\right)^{2-\frac{d}{2}}-\left(\frac{x}{\lambda^{2}}\right)^{2-\frac{d}{2}}}{D^{2}(2-d)(4-d)}-\frac{(\frac{x}{\lambda^{2}})^{1-\frac{d}{2}}t}{D(2-d)}\right]$
$=$ $\frac{R^{2}\lambda^{d-4}}{(4\pi)^{\frac{d}{2}}D^{2}(d-2)(d-4)}\left\{\lambda^{2}e^{2Dt\lambda^{2}}\left[\Gamma\left(4-\frac{d}{2},2Dt\lambda^{2}\right)-2Dt\lambda^{2}\Gamma\left(3-\frac{d}{2},2Dt\lambda^{2}\right)\right]\right$.
$\hskip 227.62204pt-\left.\left[\Gamma\left(4-\frac{d}{2}\right)+(4-d)Dt\lambda^{2}\Gamma\left(3-\frac{d}{2}\right)\right]\right\}.$
@@
In the second line above the transform $y=\frac{x}{\lambda^{2}}+2Dt$ is used in the first term and then upper incomplete gamma functions can be used. Note both factors d2d-2 and d4d-4 are removable singularities, which can be seen by L’Hopital’s rule upon the second line. For finite $t$ the incomplete upper gamma functions are entire and contain no divergences. The only divergences arise from the final two terms for dimensions $d=6,8,10,\dots$ and the term is otherwise regulated. As $t$ approaches zero, the incomplete gamma functions become complete (with poles), also presenting divergences at these dimensions. This matches the primitive dimension of divergence. Thus the UV regulation is found to be weaker than for the normal interaction, which has an exponentially decaying tail and the integral is convergent in any dimension. For $d<2$ we again find IR behaviour for large $t$.
The cross-over timescale between local and non-local features can also be observed for the screened Poisson form of interaction. Firstly, from the second line in the equation above, we again see that in the limit of $\lambda\rightarrow\infty$ the term $\frac{R^{2}t^{-\frac{d}{2}+2}}{(8\pi D)^{\frac{d}{2}}(2-d)(4-d)}$ for the local model is recovered. Furthermore, from the same line we find expansion $\frac{1}{\varepsilon}[(\frac{x}{\lambda^{2}}+2Dt)^{\frac{\varepsilon}{2}}-(\frac{x}{\lambda^{2}})^{\frac{\varepsilon}{2}}]=\log\sqrt{1+\frac{2Dt\lambda^{2}}{x}}+\mathcal{O}(\varepsilon)$ (where $\varepsilon=4-d$), demonstrating the removable singularity at dimension $d=4$ for finite $\lambda$. We see that only the leading term remains at this dimension, which diverges as $\lambda\rightarrow\infty$, as would be expected. Note that if $\lambda$ is increased, $t$ must decrease to maintain a fixed value for this term. This implies that as $\lambda$ is increased and the model becomes more localised, the range of $t$ values exhibiting effective UV regulation is reduced.
For the spherical interaction, a precise calculation is somewhat intractable, although we note the following. For general interaction profile $R(k)=R\hat{R}(k)$ we find that,
$I_{2}$ $=$ $R^{2}\int_{t>t_{2}>t_{1}>0}\fint_{k}\hat{R}(k)^{2}e^{-2Dk^{2}(t_{2}-t_{1})}=\frac{2R^{2}}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{d}{2})}\int_{0}^{\infty}\mbox{d}k\hskip 1.42262ptk^{d-1}\hat{R}(k)^{2}\left[\frac{t}{2Dk^{2}}-\frac{1}{4D^{2}k^{4}}+\frac{e^{-2Dk^{2}t}}{4D^{2}k^{4}}\right]$
$=$ $\frac{R^{2}}{(8\pi D)^{\frac{d}{2}}\Gamma(\frac{d}{2})}t^{2-\frac{d}{2}}\int_{0}^{\infty}\mbox{d}x\hskip 1.42262ptx^{\frac{d}{2}-1}\hat{R}\left(\sqrt{\frac{x}{2Dt}}\right)^{2}\frac{x-1+e^{-x}}{x^{2}}.$
Now, any form of asymptotic decay of the function $\hat{R}(\sqrt{\frac{x}{2Dt}})$ will render the integral convergent at critical dimension $d_{c}=2$. Any positive power law for $t$ from this term will also push the divergence of $I_{2}/I_{1}$ as $t\rightarrow 0$ to dimensions above $d_{c}$ and UV divergence will be regulated.
Then, for the spherical interaction, we have the asymptotic form given by $R(k)\propto(\frac{k}{\lambda})^{-\frac{d}{2}}J_{\frac{d}{2}}(\frac{k}{\lambda})\sim(\frac{k}{\lambda})^{-\frac{d}{2}}\sqrt{\frac{2\lambda}{\pi k}}\cos(\frac{k}{\lambda}-(d+1)\frac{\pi}{4})$. Note that the higher the dimension, the stronger the asymptotic decay. We subsequently find that integral $I_{2}$ is finite in any dimension and $I_{2}/I_{1}\sim\mathcal{O}(\lambda^{d+1}t^{\frac{3}{2}})$ meaning UV divergence is regulated in any dimension for small enough $t$, although we again will have IR behaviour for asymptotic large $t$.
Consider the cross-over time scales between local and non-local behaviour for this spherical interaction profile. The $\mathcal{O}(\lambda^{d+1}t^{\frac{3}{2}})$ term means that as $\lambda$ increases and the model becomes increasingly local (and this term more divergent), maintaining the UV regulation is restricted to smaller values of $t$. An exact expression for the calculation of $I_{2}$ would make the cross-over time scale here a bit more transparent.
@@
In summary, we have shown that when the precision parameter is finite, various forms of non-local interaction act as a natural form of UV regulation (in the renormalisation sense). This suggests control of UV divergences is possible provided the interaction function tails off sufficiently quickly in momentum representation. We have also seen that as the precision parameter $\lambda$ for any of these non-local models gets larger, the regulation weakens as the interaction tends toward local behaviour and UV divergences start to emerge. Even for finite precision though, IR behaviour is still present for large $t$, for which we turn to rescaling methods that underlie the renormalisation group.
@@ 3.2 Rescaling
Here it is shown that the renormalisation group calculation in [23] can also be found via a simple rescaling of the action in Eq. (4). Specifically, if we perform the following space-time-field transforms,
@@
Without renormalisation, this form of rescaling is not too involved. For example, it is a simple exercise to show the mean field solution in Eq. (7) satisfies this property. However, once renormalization is set at some scale, establishing this form of invariance is less straightforward. This expression is also somewhat akin to the solution to the Callan-Symanzyk equations seen in [23]. The main differences are firstly, there is rescaling of the extra parameter $\lambda$, and secondly, that it is in terms of the original interaction parameter $R$ rather than a renormalised version. However, we can recover the Callan-Symanzyk solution via the simple scaling above (without solving or needing the Callan-Symanzyk PDE) as follows.
In the case of local interactions, the following vertex function is obtained in momentum-Laplace space (see [23] for details),
@@
The constant $g^{*}=\frac{(4\pi)^{1-\frac{\epsilon}{2}}}{\Gamma(\frac{\epsilon}{2})}$ is of order $\mathcal{O}(\epsilon)$. This is just the expression given in [23] (which considers the more general (but local) $K$ particle process $KA\rightarrow\phi$) restricted to the case $K=2$. Next we select a fixed momentum scale $\kappa$ by defining $\kappa^{2}=s+p^{2}/2$ for some $s$ and $p$. In [23] and other applications it is convenient to pick zero momentum $p=0$ (and so $s=\kappa^{2}$), although the specific choice is immaterial and somewhat arbitrary. This is used to define a renormalised parameter $R_{r}$,
@@
This is precisely the form seen in [23] as part of the Callan-Symanzik solution. In particular, for $d<d_{c}=2$ we find that as $t\rightarrow\infty$ (so $\gamma\rightarrow 0$), the limit $g_{r}^{\prime}\rightarrow g^{*}$ arises, whereas for $d>d_{c}$ the limit $g_{r}^{\prime}\rightarrow 0$ occurs. These are precisely the fixed points of the Callan-Symanzyk equations in [23]. Note also that the relationship between $g_{r}$ and $g_{r}^{\prime}$ as we shift between $t$ and $t^{\prime}$ is independent of the choice $\kappa$, which does nothing other than set the scale and initialise the value $g_{r}$. Finally, we note that if the $R$ dependency is replaced with the dimensionless parameter $g_{r}$ (and we write $g_{r}^{\prime}=g_{r}(\gamma)$), the rescaling law in [23] is found, that is, the full solution to the Callan-Symanzyk equation for the density;
@@
Note that from $\gamma^{2}=\frac{t^{\prime}}{t}$ we find that for large time, the precision parameter, scaling as $\lambda\gamma^{-1}$, becomes large, and the interaction profiles become local. This ultimately explains why the asymptotics of non-local systems are the same as local systems, although the cross-over time scale of such behaviours is a natural question to consider, as briefly explored in the previous section.
@@ 3.3 The Response Functional
The response functional is the standard propagator running from time $t_{1}$ to $t_{2}$ with any number of tree branches attached, where each tree runs from time 0 to the time of attachment somewhere in $t_{1},t_{2}]$. The first few corresponding diagrams can be seen in Fig. 5 (represented as spirals). The external momenta running along attached trees are taken to be zero (this essentially means the initial densities $n_{0}$ at the end of trees are spatially uniform). The momenta running along the response functional, $k$ (which may be part of a loop in a larger diagram, for example), can be non-zero. Note that each each time point of attachment, we have two possible nodes to attach the tree branch to. This results in two possible choices for the interaction propagator, either $R(0)$ or $R(k)$, leading to the expression,
@@
$=e^{-Dk^{2}(t_{2}-t_{1})}\left(\frac{1+Rn_{0}t_{1}}{1+Rn_{0}t_{2}}\right)^{\frac{R(0)+R(k)}{R}}=e^{-Dk^{2}(t_{2}-t_{1})}\left(\frac{1+Rn_{0}t_{1}}{1+Rn_{0}t_{2}}\right)^{1+\hat{R}(k)}.$